OpenStudy (anonymous):

The sequence is geometric. What is the next term? 3125, 625, 125, 25, ....

OpenStudy (turingtest):

might want to check to see if each term has a common ratio...

OpenStudy (anonymous):

\[\sum_{?}^{?}a^n\]

OpenStudy (anonymous):

well backwards if you take your sequence and flip it

OpenStudy (anonymous):

\[\sum_{?}^{?}5^n\]

OpenStudy (turingtest):

\[\frac{a_{n+1}}{a_n}=\frac{625}{3125}=\frac{125}{625}=\frac{25}{125}=R\]once you know the common ratio R just multiply to find the next term\[a_{n+1}=Ra_n\]